# ChOiRe: Characterizing and Predicting Human Opinions with Chain of Opinion Reasoning

### v3

@@ 2 Related Work
Large-scale model architectures Devlin et al. (2019); Radford et al. (2019); Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023) have enabled large language models (LLMs) to excel at various NLP tasks using zero- or few-shot prompting Liu et al. (2023). Notably, Wei et al. (2022); Kojima et al. (2022) propose prominent Chain-of-Thought (CoT) techniques, enabling LLMs to explicate intermediate reasoning steps to solve multi-step reasoning tasks with higher fidelity and efficiency.
Can CoT analyze and predict human opinion effectively? We find that a naïve application of CoT does not help GPT-X models (section 5), but that an appropriate modification does. We propose Chain-of-Opinion (CoO) reasoning (section 3) that overcomes CoT’s limitations in this task. Noting that prompting techniques such as task decomposition Khot et al. (2023); Zhou et al. (2023) and retrieved-based methods Yao et al. (2023); Shinn et al. (2023) have been recently introduced, we focus only on the reasoning explanation aspect here given the abstractive and challenging nature of the task.
Figure 1: ChOiRe overview, consisting of the four main steps (cyan background), as detailed in section 3.
3 ChOiRe: A Chain of Opinion Framework
Task Formalisation. We follow Santurkar et al. (2023), and formulate the opinion prediction task as multiple-choice question answering. Formally, a benchmark with $N$ data points is denotated as $D=\{\langle T,E,I,q,a\rangle_{n}\}_{n=1}^{N}$, where $T$, $E$ and $I$ indicate the (T)opic of a question $q$, the (E)xplicit personae and (I)mplicit personae of the user answering $q$ with opinion $a$. Following the prior work, $E$ consists of $12$ user demographic and ideology metadata attributes, and $I$ contains a number of the user’s historical opinions in the format of question–answer pairs. Models then learn to analyze $T,E,I,q$ and predict the opinion $a$.
Fig. 1 shows an overview of ChOiRe, consisting of four main steps (marked with a cyan background). First, ChOiRe employs an LLM to analyze and select a subset of relevant explicit personae, denoted as $E^{rel}\subseteq E$ for answering the opinion question $q$. The LLM then assesses the informativeness of the implicit personae ($I$) in predicting $q$, selecting the top-$K$ implicit personae (termed LLMtop-$K$). Next, an LLM is prompted to explain the provided explicit $E^{rel}$ and implicit LLMtop-$K$ personae sequentially in a Chain-of-Opinion (CoO) reasoning strategy. Finally, ChOiRe calls the LLM to predict the opinion $a$ with varying values of $K$ for the top-$K$ implicit personae. ChOiRe chooses the opinion with the highest frequency as the final prediction We detail the steps below.
3.1 Filtering Explicit Personae Attributes (FEA)
Accounting for explicit personae, which consist of the demographic and ideological metadata attributes of users — such as their age, income, and political ideology — has been shown to help models characterize and predict human opinions more accurately Hwang et al. (2023). However, which personae matter and which do not? are still open questions. Section E.1 shows such an example in full, where, when considering all of explicit personae, the model makes an incorrect prediction while removing unnecessary personae the model made a correct prediction. This may be caused by the LLM’s attention mechanism’s forcing the model to attend to all input tokens, even irrelevant ones. To filter out unnecessary explicit personae, we ask the LLM to reason and analyze how each persona is helpful for the model to predict the opinion via Chain-of-Thought to output a list of which personae are relevant given the question and the opinion answer choices [^4] We provide our FEA prompt in Section C.1. Surprisingly, we find that LLMs evaluate more than half of the explicit personae as irrelevant on average. We further conduct human evaluations to verify this finding in section 4.
3.2 Implicit Personae Opinions Ranking (LLMtop-$K$)
LLMs have been sensitive to selected demonstrations and their order in the prompts Perez et al. (2021); Luo et al. (2023); Gao et al. (2023). For predicting human opinions, we discover that LLMs are also sensitive to the chosen implicit personae opinions as input. Hwang et al. (2023) rank the implicit personae opinions via semantic-similarity scores and select top-$K$. This strategy is suboptimal because the top-ranked opinions in terms of semantic similarity may not be the ones that provide the most supportive information for the models to predict opinions (section E.2). As LLMs are shown to be good data analysts Wang et al. (2023a); Cheng et al. (2023), we propose to address the above challenge by utilizing LLMs to analyze and rank the implicit personae opinions in usefulness descending order. Our finding is that despite the output rankings from LLMs varying with different input orders of implicit personae opinions, the sets of LLMtop-$K$ opinions overlap by a good coefficient when $K$ is large enough ($\geq 8$) (fig. 5). Therefore, we propose to input the implicit personae opinions to LLMs in a random order to make our method more versatile. We also examine the case where we input the opinions in the semantic similarity order. We illustrate the prompt template in section C.2. By performing this step, our proposed method supports the usefulness of opinions in predicting the test opinions, rather than the semantic similarity. We term this method as LLMtop-$K$.
3.3 Chain-of-Opinion Reasoning (CoO)
Wei et al. (2022); Kojima et al. (2022) introduce few-shot and zero-shot Chain-of-Thought (CoT) prompting strategies demonstrating that by reasoning step-by-step, LLMs can achieve promising results on complex tasks. However, the sampled reasoning steps can be inconsistent, leading to possibly different outcomes Wang et al. (2023b). Furthermore, it is little known how the models perceive multiple implicit personae opinions, especially when many opinions are input, which one(s) the models used, which one(s) they didn’t for predicting the opinion? Our preliminary experiments with CoT (sections 6.1, E.4 and E.3) reveal that the CoT explanations can vary frequently based on different subsets of opinions mentioned in their explanations, leading to diverse final answers, especially when the decoding temperature is relatively high555greater or equal to 0.6. To mitigate this issue, we propose to instruct the LLMs to analyze the given explicit and implicit personae one by one before concluding the prediction via simply adding "explaining and analyzing how each of the Opinions and Demographic Information supports the question" into the prompt instruction. Given an LLM that can follow human instructions well such as ChatGPT OpenAI (2022), this addition offers two notable advantages. First, for each question, we ensure that the model explains and analyzes the provided personae one by one without missing any, possibly resulting in more thorough predictions. Second, this method helps the model to output more consistent reasoning explanations, enhancing its reliability (section 6.1).
3.4 Answer Consistency with Dynamic Numbers of Opinions
Prior work Hwang et al. (2023) fixes the number of implicit personae opinions for prediction to $K=8$. However, this approach occasionally results in models generating "...the answer cannot be determined." (tables 4 and E.5). We attribute this to insufficient user implicit personae opinions provided. Inspired by Self-Consistency (SC) Wang et al. (2023b), our approach involves sampling multiple answers using different $K$ values for a given question. The most frequent answer, along with its explanation, becomes the final prediction. Our method is distinct from SC since SC samples multiple answers with a fixed prompt. We experiment with $K\in\{8,10,12\}$ for efficiency.
4 Evaluation
Dataset.
We experiment on OpinionQA dataset Santurkar et al. (2023) — the only opinion QA dataset to date consisting of both user explicit and implicit personae designed for the assessment of alignment between LMMs’ opinions and human participants, encompassing a diverse range of $60$ US demographic groups. It covers $60$ US demographic groups, with $15$ topics, each comprising around $100$ questions, gathered from $5,340$ users.
Dataset Preprocessing.
Due to limited resources, we randomly sample $25$ users per topic for our experiments. For each user, we follow Hwang et al. (2023) to use $20\%$ of the implicit questions as the implicit persona. For the remaining $80\%$ implicit questions, we randomly select a maximum of $15$ implicit questions for testing. Our sampling method results in a total of $375$ users and $5,603$ implicit evaluation question–answer pairs. Our subset is highly representative because we gather 25 users from every topic and 15 questions per user. Rigorous statistical tests further validate the significance of our results which align closely with Hwang et al. (2023) testing on a larger subset using InstructGPT.
Prompting Baselines.
We use both closed-source ChatGPT OpenAI (2022), ChatGPT-Instruct OpenAI (2023a), GPT-4 OpenAI (2023b), and open-source Mistral-7B-Instruct-v.02 Jiang et al. (2023) as our LLMs, and compare ChOiRe with $5$ prompting methods: (1) W/o persona, where LLMs are evaluated without user historical opinions, ideology, or demographic data; (2) Demographic + Ideology + top8 Opinions (termed DIO-top8), introduced by Hwang et al. (2023) demonstrating that integrating explicit and implicit personae enhances user opinion modeling and prediction, achieving state-of-the-art results on OpinionQA at that time; (3) DIO-top8 + CoT is the Chain-of-Thought (CoT) prompting Kojima et al. (2022) version of DIO-top8 involving appending "answer the following question step-by-step" to prompts, aiming to explore whether CoT improves model performance in this task; (4) DIO-top8 + SC is the baseline which we apply the Self-Consistency technique with CoT Wang et al. (2023b) to DIO-top8 to select the most frequent answer generated by the model as the final opinion prediction; (5) DIO-top8 + Self-refine Madaan et al. (2023) interactively feedbacks and refines the answers by LLMs. We do not experiment with InstructGPT Ouyang et al. (2022) like Hwang et al. (2023) since this model is going to be deprecated and replaced by ChatGPT-Instruct. For GPT-4, we only run the main experiment and we use ChatGPT for FEA and LLMtop-$K$ steps due to our limited budget. All the prompts and costs are in appendix C, implementation details in section A.1, and more baselines in section B.1.
Fine-tuning Baselines.
We further investigate whether ChOiRe’s FEA and LLMtop-$K$ steps (section 3) also improve fine-tuning for opinion-aligned models. We first create the fine-tuning data by using ChatGPT to perform ChOiRe’s FEA and LLMtop-$K$ steps ($K=8$) on a training set of $30,000$ samples randomly selected from OpinionQA which are different from our $5,603$ test ones. We then fine-tune and evaluate GPT-2 models (base, large) Radford et al. (2019) and FlanT5 models (base, large) Chung et al. (2022). Fine-tuning details are provided in section A.1.
Metrics.
We employ Accuracy and Collapsed Accuracy as the automatic evaluation metrics following Hwang et al. (2023). It is worth noting that Precision/Recall/F1 is not applicable in our task, since the numbers of answer choices are not the same for all the OpinionQA samples. In addition, human evaluations are crucial due to the absence of automated metrics assessing LLMs’ performance in FEA, LLMtop-$K$ and CoO steps of ChOiRe. Therefore, we conduct our human assessments to address these research questions: (1) LLMs’ effectiveness in filtering unnecessary explicit personae; (2) LLMs’ proficiency in ranking implicit personae opinions; (3) LLMs’ ability to explain answers via CoO. To this end, we randomly select $100$ answers generated by ChOiRe with ChatGPT, ChatGPT-Instruct, GPT-4, Mistral. We then hire $3$ excellent undergraduates who are native English speakers as annotators. For FEA and LLMtop-$K$ steps, each annotator is instructed to rate on a 1-3 scale (3 is best) via the Satisfaction criterion defined as how well the algorithm of LLMs performs in filtering/ranking, subjectively. To answer (3), we use two criteria named Reasonableness measuring how well the LLMs reason with the CoO explanations, and Follow the Instruction assessing the capability of LLMs in following our instruction to explain and predict the opinions. Three annotators are also guided to rate the criteria on a 1-3 scale. Each metric’s final score is the average of three annotators’ scores. The scoring instructions are in section D.1 and the inter-annotators’ agreement is assessed by Kripp’s alpha Krippendorff (2011).
5 Main Results
Model ChatGPT ChatGPT-Inst GPT-4 Mistral-7B-Ins.-v0.2
W/o persona 46.60/65.72 44.91/63.60 - 41.24/59.54
DIO-top8 50.22/69.21 51.95/71.16 57.98/76.86 44.16/62.47
DIO-top8 + Self-refine 43.14/65.33 42.71/62.98 - 36.23/55.06
DIO-top8 + CoT 49.96/69.05 51.90/71.51 - 52.25/71.95
DIO-top8 + SC 50.58/69.66 52.06/71.87 - 53.14/72.88
DIO-top8 + FEA 50.64/69.85 52.63/72.30 - 44.99/64.09
DIO-top8 + CoO 50.97/70.22 52.08/71.65 - 53.79/73.59
DIO-LLMtop8 51.03/70.31 52.80/72.60 - 45.86/64.98
DIO-LLMtop8 + FEA 51.19/70.69 52.97/72.84 - 45.23/64.73
DIO-LLMtop8 + FEA + CoO 51.90/71.57 53.01/72.91 59.02/78.70 54.21/74.09
ChOiRe 52.21normal-$\dagger$/72.09normal-$\dagger$ 53.26normal-$\dagger$/73.26normal-$\dagger$ 59.30normal-$\dagger$/78.82normal-$\dagger$ 54.43normal-$\dagger$/74.34normal-$\dagger$
% Improvements +3.22/+3.49 +2.52/+1.93 +2.28/+2.55 +2.42/+2.00
Table 1: Overall Accuracy/Collapsed Accuracy on ChatGPT, ChatGPT-Instruct, and GPT-4. FEA is our first step, Filtering Explicit Attributes. LLMtop8 is the second step, ranking and selecting top-8 implicit persona opinions as demonstrations, CoO is Chain-of-Opinion reasoning. Improvements are calculated with the best baseline. $\dagger$ denotes our model outperforms baselines significantly with p-value < 0.01 under t-test.
Model GPT-2-base GPT-2-large FlanT5-base FlanT5-large
W/o persona 41.14/58.87 21.94/39.11 48.98/68.33 39.83/58.43
DIO-top8 21.23/38.64 24.94/42.22 55.00/74.98 54.94/74.79
DIO-top8 + FEA 22.62/40.97 25.65/45.21 55.78/75.34 58.77/77.26
DIO-LLMtop8 22.65/41.12 28.86/47.60 57.97normal-$\dagger$/77.46normal-$\dagger$ 58.20/77.56
DIO-LLMtop8 + FEA 25.05/44.41 29.54normal-$\dagger$/48.66normal-$\dagger$ 57.45/77.13 59.00normal-$\dagger$/78.46normal-$\dagger$
% Imp. over DIO-top8 +17.99/+14.93 +18.44/+15.25 +5.40/+3.30 +7.38/+4.90
Table 2: Performance of fine-tuned baselines with our proposed FEA and LLMtop8 steps preprocessed by ChatGPT. $\dagger$ denotes our model significantly outperforms baselines with p-value < 0.01 under t-test.
Overall Prompting Results.
Table 1 shows our macro experimental outcomes. We derive $4$ main observations in this task. First, ChOiRe improves the best among baselines significantly with $3.22\%$, $2.52\%$, $2.28\%$, $2.42\%$ accuracy for ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral. It establishes a strong SOTA result with GPT-4, surpassing previous SOTA DIO-top8 with InstructGPT achieving 53.74% Hwang et al. (2023) by a notable margin. Notably, in the case of GPT-4, we utilize ChatGPT for FEA and LLMtop-$K$ steps, showcasing the strength of a weaker model that enhances a stronger one. Second, we see that Accuracy and Collapsed Accuracy have the same trend, and ChOiRe also achieves the SOTA on Collapsed Accuracy with the highest improvement of $3.94\%$ observed with ChatGPT. Third, naïve CoT ("answer the following question step-by-step") helps Mistral but slightly harms ChatGPT and ChatGPT-Instruct with DIO-top8 Hwang et al. (2023). On the other hand, SC improves all models. Therefore, we attribute CoT’s limitation to the inconsistency of its explanations (3). Meanwhile, ChOiRe with CoO consistently attains improvements, verifying the effectiveness of explicitly requiring the model to analyze all the personae. Finally, ChatGPT, ChatGPT-Instruct, and Mistral show improvements by selecting only $4.79/12$ and $5.59/12$, $8.83/12$ explicit personae on average, respectively. This suggests that over half of explicit personae may be noisy for models to predict opinions.
Fine-grained Prompting Results.
Diving deeper into the benchmark topics in table 5, ChOiRe achieves SOTA results in $8/15$, $8/15$, $11/15$, $13/15$ topics for ChatGPT and ChatGPT-Instruct, GPT-4, and Mistral. The improvements are especially huge for some topics. For example, compared with the best among baselines, it improves GPT-4 up to $12.08\%$ accuracy on Views on gender, ChatGPT up to $9.82\%$ on Economic Inequality. We also specifically compare ChOiRe with the best baseline DIO-top8 + SC in section B.3, showing $8/12$ improvements for ChatGPT and ChatGPT-Instruct. We further plot the accuracy distribution over users of ChOiRe, specifically for ChatGPT in fig. 4. We see that the majority accuracy is $0.5$, with a few users scoring zero and over $20$ achieving perfection.
Fine-tuning Results.
Table 2 presents our fine-tuning outcomes. Notably, leveraging the ChOiRe’s FEA and LLMtop-$K$ steps on the fine-tuning data yields substantial enhancements for GPT-2-large and FlanT5-large, showcasing relative accuracy improvements of $18.44\%$ and $7.38\%$ respectively. Remarkably, ChOiRe’s FEA and LLMtop-$K$ steps bring FlanT5-large’s performance on par with GPT-4, despite GPT-4’s significantly stronger capability. Furthermore, ChOiRe’s LLMtop-$K$ proves particularly beneficial for enhancing FlanT5-base. Surprisingly, GPT-2-base performs well even without user demographic and ideological information, possibly due to potential contamination Sainz et al. (2023) with public polling data from OpinionQA.
Human Evaluation Results.
Model FEA Satis. LLMtopK Satis. Rea. Foll. Inst.

### v5

@@ 2 Related Work
Large-scale model architectures Devlin et al. (2019); Radford et al. (2019); Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023) have enabled LLMs to excel at various reasoning NLP tasks via prompting Wei et al. (2022); Khot et al. (2023); Zhou et al. (2023); Wang et al. (2023b); Shinn et al. (2023). Notably, Wei et al. (2022); Kojima et al. (2022) proposed the popular Chain-of-Thought (CoT) techniques, enabling LLMs to explicate intermediate reasoning steps, aiding the solving of multi-step reasoning tasks with higher fidelity and efficiency.
Can CoT analyze and predict human opinion effectively? Surprisingly, a naive application of CoT fails to improve GPT-X models (Section 5.2). We attribute this to the reasoning inconsistencies of CoT and the complexity of the task: strategic reasoning is essential to consistently fully utilize the nuanced explicit and implicit personae. Our VBN reasoning (Section 4) overcomes CoT’s limitations.
3 LLMs are Distracted by Irrelevant Personae
Type Relevant +1 less/irr +3 less/irr +all less/irr
persona relevant relevant relevant
ChatGPT Explicit 37.56 35.53$\downarrow$ 34.51$\downarrow$ 34.35$\downarrow$
Implicit 34.35 33.84$\downarrow$ 30.76$\downarrow$ 31.28$\downarrow$
LLaMa 3.1 Explicit 26.39 24.36$\downarrow$ 23.85$\downarrow$ 22.84$\downarrow$
Implicit 26.66 25.12$\downarrow$ 21.02$\downarrow$ 23.08$\downarrow$
Table 1: The performances of ChatGPT (gpt-3.5-turbo-0125) and LLaMa3 (8B 3.1-it) significant drop (approx. 1-6%) on Gun topic of OpinionQA Santurkar et al. (2023) when less or irrrelevant personae added.
We study the sensitivity of LLMs to irrelevant personae which motivates COO in Section 4. This exploration is related to Shi et al. (2023) but we quantify the LLM sensitivity to personae instead. We find that LLMs can be easily distracted by explicit or implicit personae that are less or irrelevant to opinion questions. To examine this, we perform a semi-human evaluation to assess the relevancy of personae on 197 randomly chosen Guns questions from the OpinionQA dataset Santurkar et al. (2023). Each sample is denoted as $\{T,E,I,q,o,a\}$ where $T$, $E$, and $I$ indicate the topic, explicit personae (demographics and ideology), and implicit personae (historical opinions) of the user answering $q$ with opinion options $o$ and correct label $a$.
To assess the relevance of each explicit persona in $E$ to $q$, we employ two native English undergraduates and ChatGPT (gpt-3.5-turbo-0125) OpenAI (2022). The annotators carefully examined ${q,a}$ and labeled 12 attributes of $E$ as relevant or irrelevant, determined by majority vote. The annotators achieve a good agreement of 60.2% Krippendorff’s alpha Krippendorff (2011). For implicit personae ($n=\sim 20$), assessing their relevance to $q$ by manual means is costly. Therefore, we compute the semantic similarity between them and $q$ using OpenAI text-embedding-ada-002 and label the top-8 as relevant and the rest as irrelevant.
We test four different setups for each type of personae: (i) using only relevant ones; (ii) including one, (iii) three, or (iv) all irrelevant ones. To ensure that our results are consistent, we use Self-Consistency Wang et al. (2023b) with $5$ times sampling. We experiment with both representative closed and open source LLMs: ChatGPT and LLaMa3 Touvron et al. (2023).
Results in Tab. 1.
Surprisingly, adding a single irrelevant explicit persona results in a prediction change of $30\%$ for ChatGPT and $40\%$ for LLaMa, with $2\%$ performance drop for both. Meanwhile, adding one irrelevant implicit causes much smaller drops for both. Explaining this phenomenon, we have two hypotheses: first, both models rely more on explicit personae for opinion prediction; second, the so-called irrelevant implicit persona may still hold some relevance, as low semantic similarity does not entirely equate to low relevance (Appx. Fig. 6). Additionally, we observe that adding three/all irrelevant for both persona types significantly reduces performance by over 3-5% in absolute, indicating that irrelevant personae harm model predictions significantly.
4 COO: A Chain of Opinion Framework
These above findings suggest that for pre-trained LLMs, the choice of (relevant) personae as input is important and significantly impacts the model’s outcomes. Notably, we observe that training language models sees even greater benefits when incorporating only relevant personae. We introduce a simple and cost-efficient framework, termed Chain-of-Opinion (COO) characterizing explicit and implicit personae for opinion prediction. COO serves as an intermediate data preprocessing step for fine-tuning and prompting. We assume optional access to the user’s explicit and implicit personae. Let ${G}_{\mathcal{M}}:\mathcal{V}^{*}\to\mathcal{V}^{*}$ be the generation function of $\mathcal{M}$ where $\mathcal{V}$ is the model vocabulary, $Q$ be the concatenation of the test question and its answer choices, and $a$ be the correct label.
4.1 COO in Prompting
As introduced in Section 1 and depicted in Fig. 1, COO distinctly processes explicit and implicit personae, motivated by the Value–Belief–Norm theory Stern et al. (1999), and consists of four main steps: Step 1. Filtering explicit personae; Step 2. Ranking implicit personae; Step 3. Value-Belief-Norm (VBN) reasoning; and Step 4. Answer consistency with dynamic numbers of opinions.
Figure 1: COO overview with four main steps marked with the nuts. It processes explicit and implicit personae parallelly to facilitate the missing personae scenarios.
Step 1. Filtering explicit personae (FEA).
This step aims to binarily filter explicit personae (the set $E$, denoted in Section 3), including user demographics and ideology, for prediction. Irrelevant ones can harm the model performance significantly (Section 3), possibly because of LLM’s attention mechanism forcing the model to attend to all input tokens, including irrelevant ones. Since the relevancy of an explicit persona to a test question is apparent, we binarily filter out irrelevant explicit personae by instructing the LLM $\mathcal{M}$ to reason and analyze how each of them is helpful for the model to predict the opinion via Chain-of-Thought Wei et al. (2022):
$E_{rel}:=G_{\mathcal{M}}([E,Q])$ (1)
The instruction for $\mathcal{M}$ is skipped in Eq. 1. Surprisingly, LLMs evaluate more than half of the explicit personae as irrelevant on average. We conduct human evaluations to examine this in section 5, and provide an example in full in Section E.2 showing when considering all explicit personae, the model yields an incorrect prediction while removing unnecessary personae, it offers a correct one.
Step 2. Implicit personae opinions ranking (LLMtop-$K$).
This step focuses on ranking and selecting implicit personae ($I$) consisting of user historical opinions for prediction. Identifying the most supportive ones for a test question is critical yet complex and often requires significant comprehension efforts. Hwang et al. (2023) utilized semantic-similarity between ($q$, $i$) to select implicit persona $\forall i\in I$. This strategy is suboptimal because the top semantic similarity opinions may not be the ones that provide the most supportive information for the models (Section E.2, also we hypothesized in Section 3). As LLMs are shown to be good data analysts Wang et al. (2023a); Cheng et al. (2023), we propose to utilize LLMs to analyze and rank the implicit personae opinions in usefulness order:
$I_{rel}:=top_{K}(G_{\mathcal{M}}([I,Q]))$ (2)
The instruction of $\mathcal{M}$‘ is skipped in Eq. 2. We fix $K$ when selecting top-$K$ implicit personae rather than having LLMs directly select useful ones like COO’s Step 1 because implicit personae involve more complex and nuanced information with varying degrees of relevance, unlike the binary relevance of explicit personae (Section 1). Additionally, we input $I$ to LLMs in a random order to enhance the versatility of our method. We validate this approach in Section 6. Our method prioritizes supportiveness in reasoning about opinions, unlike conventional demonstration selection focusing on semantic similarity or diversity Xu et al. (2024c).
Step 3. Value-Belief-Norm reasoning (VBN).
How can we best utilize $E_{rel}$ and/or $I_{rel}$ for opinion prediction? Popular prompting methods for reasoning, such as few-shot and zero-shot Chain-of-Thought (CoT) Wei et al. (2022); Kojima et al. (2022), typically instruct LLMs to reason step-by-step. However, for opinion prediction, a general step-by-step guide presents two key issues. First, the generated reasoning steps can be inconsistent, causing divergent outcomes, particularly at medium-high decoding temperatures ($\geq$ 0.6) (proven in Section 6.2 and Section E.2). This significantly undermines the reliability and accuracy of models. Second, step-by-step reasoning lacks interpretability, as it remains unclear how models process and prioritize multiple explicit and implicit personae. When numerous opinions and personae are input, it is difficult to discern which were used in forming the prediction and which were disregarded.